OpenStudy (anonymous):

Pls help divide and simplify. 9√(7) / 45√(35)

OpenStudy (anonymous):

\[\frac{9\sqrt{7}}{45\sqrt{35}}\]correct?

OpenStudy (anonymous):

Divided by in btwn

OpenStudy (anonymous):

Didn't show up at first but yes yours is correct

OpenStudy (anonymous):

I git either1/Sqrt5 or sqrt5/25

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

frits divi 9 top and bottom \[\frac{\sqrt{7}}{5\sqrt{35}} \] \[\frac{\sqrt{7}}{5\sqrt{35}}\times\frac{5\sqrt{35}}{5\sqrt{35}}\]

OpenStudy (anonymous):

\[\frac{5\sqrt{245}}{25\times35} =\frac{5\times7\sqrt{5}}{875}\]

OpenStudy (anonymous):

\[\frac{35\sqrt{5}}{875}=\frac{\sqrt{5}}{25}\] here answer

OpenStudy (anonymous):

thank